OpenStudy (anonymous):

Solve. 1/25(<---this is a fraction) = 5x + 4

OpenStudy (anonymous):

sub. 4 to both sides 1/25-4= -3.96 -3.96=5x div. 5 to both sides x=0.792

OpenStudy (anonymous):

srry x= -0.792

OpenStudy (anonymous):

hold on a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/25=5x+4 -4 -4 -3.96= -3 24/25 div. 5 div. 5 x=-99/125

OpenStudy (anonymous):

when you want to simplify a decimal you put it over a hundred and simplify it. that's only if the decimal is 2 places. if its three its over a thousand..... so forth

OpenStudy (anonymous):

gah thank u so much

OpenStudy (anonymous):

You can also get rid of fractions early on by multiplying all terms by 25: